5. 15 points Find a branch of log(z2 + iz – 3) such that it is analytic at z = i, and find its derivative at z =i.
Answer #1

Solution Find a branch of log (mz² + d2-3) 2²12-3=0 Nowe - 2 2 1 3 is Branch point of log (2²12-3) Now f(x) = log (2²+62-3) fi hope it will help you...

For any quarry comment please...